Exercise 5.85 Part A A scuba diver creates a spherical bubble with a radius of 2.5 cm at a depth of 30.0 m where the total pressure (including atmospheric pressure) is 4.00 atm What is the radius of the bubble when it reaches the surface of the water? (Assume atmospheric pressure to be 1.00 atm and the temperature to be 298 K.) Express your answer using two significant figures. r= cm Submit My Answers Give Up Provide Feedback Continue
Answer #1

P1 = 4.00 atm
V1 = 4/3*pi*r1^3
P2 = 1.00 atm
V2 = 4/3*pi*r2^3

Use Boyle’s law:
P1*V1 = P2*V2
4.00*4/3*pi*r1^3 = 1.00*4/3*pi*r2^3
4.00*r1^3 = 1.00*r2^3
4.00*(2.5)^3 = 1.00*r2^3
r2^3 = 62.5
r2 = 4.0 m

Answer: 4.0 m